Explain what the value of the discriminate means to the graph of y=ax6+bx+c. Choose values of a,b,c to create a particular discriminate. Graph your corresponding equations.
Answer by rapaljer(4667)

Is it possible that you mean y=ax^2+bx+c? What you stated y=ax6+bx+c just doesn't make sense to me.

R^2 at SCC